# Generic rigidity and accidental modes in metal-organic frameworks

### v1

@@ IV Discussion
Our results reveal that most MOFs occupy a mechanically marginal regime. Although the majority of frameworks are formally over-constrained, their node and linker geometries introduce accidental geometric modes that make some constraints effectively dependent. These additional zero modes either destabilize the MOF’s structure or introduce soft cubic or quartic modes. Only a smaller subset of materials, such as ABIXOZ, exhibits fully independent constraints and behaves as a genuinely rigid and topologically stable network.
UiO–66, which is widely regarded as a stable MOF [UiO-66-stable, UiO-66-stable2], illustrates how geometric arrangement rather than insufficient coordination can produce internal soft modes. The accidental modes identified in the baseline model arise from specific alignments within the Zr–O clusters and linker geometries that render certain constraints redundant. These motions correspond to internal flexibilities relevant to activation and guest molecule accommodation, yet they do not compromise the overall stability of the framework. When auxiliary long-range constraints are introduced by adding bonds that arise when decreasing $\tau$, these accidental dependencies are removed, and the zero modes are lifted into low-frequency vibrational branches. This confirms that the soft motions originate from geometric degeneracy rather than from an actual deficit of constraints.
While the present rigidity-based analysis provides these mechanistic insights, it is subject to several approximations. The harmonic model captures only small amplitude motions, and quantitative accuracy depends on the UFF4MOF parameterization. The omission of nonbonded forces and entropic contributions limits the description of large deformations and finite temperature effects. Despite these simplifying assumptions, the correspondence with first-principles calculations for benchmark systems indicates that the model captures the essential degree of freedom structure that governs phonon modes.
@@
Finally, the prevalence of near-isostatic frameworks revealed here points toward a broader connection between MOF mechanics and ideas from topological mechanics. The parallels to isostatic lattices explored by Kane and Lubensky [Kane2013] suggest that similar topological features, including protected boundary modes, may be realizable in complex porous crystals.
@@ Appendix A Fourier Transformed Rigidity Matrix
$\mathbf{e}(\mathbf{k})=\mathbf{R}(\mathbf{k})\,\mathbf{u}(\mathbf{k}).$ (10)
@@
$V=\tfrac{1}{2}\,\mathbf{u}^{\mathrm{T}}R^{\mathrm{T}}KR\mathbf{u},$ (11)
@@ Appendix B Code and Data Availability
All computational scripts and input files used in this work are openly available at:
Structures with CIF sizes above 35 kB were excluded to avoid memory bottlenecks. Neighbor environments were identified using CrystalNN[crystalNN] with a search cutoff of 25 Å. For the main dataset we used the default, chemically aligned settings with x_diff_weight = 1.5, which produce coordination numbers consistent with known metal and linker environments. For the UiO–66 tuning in Fig. 4, we instead reran CrystalNN with chemical heuristics effectively turned off (distance_cutoffs = None, weighted_cn = True, x_diff_weight = 0) and then retained only neighbors with $\mathrm{weight}\geq\tau$. In this case the network is controlled purely by geometry through the $\tau$ filter: bonds and angles present in both the default and $\tau$ filtered networks are treated as primary UFF4MOF constraints, while those that appear only in the $\tau$ filtered network are interpreted as auxiliary long range constraints and assigned uniform spring constants (see the repository for specific values). The Maxwell–Calladine index was consistently evaluated at the $\Gamma$-point, while phonon dispersions were sampled along Setyawan Curtarolo [Setyawan_2010] high-symmetry paths using 100 evenly spaced points per segment.
@@ Appendix C Silicon Benchmark
Since our present implementation builds the rigidity matrix from an explicit real-space reference cell containing a complete set of distinct bond and angle constraints, we compute the phonon dispersion of silicon using the standard eight atom conventional cubic cell rather than the two atom primitive cell. The comparison in Fig. 6 highlights several important features. The overall frequency range predicted by the rigidity model is higher than in the DFT reference calculation, reflecting the stiffer effective force constants in the harmonic spring network. Along the $\Gamma\rightarrow X$ path, however, the dispersions are in close agreement, with comparable band shapes and similar degeneracies at the high-symmetry points. Along the $X\rightarrow K$ path the two spectra begin to diverge, yet the qualitative structure is still preserved. The most noticeable differences appear along the $\Gamma\rightarrow L$ path, where the relative ordering and spacing of branches do not fully align with the DFT results. Despite these discrepancies, the simplified rigidity-based model reproduces the principal features of the silicon phonon dispersion with reasonable accuracy. This agreement supports the use of the approach for capturing the dominant degree of freedom structure in more complex materials.

### v2

@@ IV Discussion
Our results reveal that many MOFs occupy a mechanically marginal regime in a constraint-based description. Although the majority of frameworks are formally over-constrained at, their node and linker geometries can introduce accidental geometric modes that make subsets of constraints effectively dependent. The resulting excess zero modes are not necessarily distributed uniformly throughout the framework; instead, Figs. 3 and 4 show that marginality is often highly heterogeneous and can be concentrated in specific substructures. In this sense, global Maxwell–Calladine counts are informative but incomplete: frameworks with similar global $\nu$ can exhibit qualitatively different low-frequency dynamics depending on where constraint redundancy occurs and which atoms participate in the soft subspace. Only a smaller subset of materials, exemplified by ABIXOZ, exhibits fully independent constraints and behaves as a genuinely rigid network for which global counting reliably predicts mechanical response.
The representative cases further show that the character of the zero-mode subspace matters as much as its size. In IKEBUV01, the zero-like motion is overwhelmingly dominated by hydrogen, indicating that the soft subspace is largely confined to linker-associated, peripheral degrees of freedom. By contrast, UiO-66 exhibits a large number of accidental zero modes while remaining dominated by hydrogen but with a substantial carbon component and only minor oxygen participation, consistent with soft motion that extends further into linker environments even while coupling weakly to the metal–oxygen node. This distinction helps clarify why frameworks that are similarly marginal by global indices can nonetheless differ in how mechanically consequential their soft modes are, depending on whether the zero-mode subspace remains terminal/peripheral or involves heavier framework atoms.
UiO-66, which is widely regarded as a mechanically stable MOF [28, 29], it is not simply rigid. To understand how the zero modes we identify in our idealized model can produce stability when going beyond our idealizations, we have to understand the limitations of our approach.
While the present rigidity-based analysis provides these mechanistic insights, it is subject to several approximations. The harmonic model captures only small-amplitude motions, and quantitative accuracy depends on the UFF4MOF parameterization. The omission of nonbonded forces and entropic contributions limits the description of large deformations and finite-temperature effects. Despite these simplifying assumptions, the correspondence with first-principles calculations for benchmark systems indicates that the model captures the essential degree-of-freedom structure that governs phonon modes.
@@
Finally, the prevalence of near-isostatic frameworks revealed here suggests a broader connection between MOF mechanics and ideas from topological mechanics. The parallels to isostatic lattices explored by Kane and Lubensky [13] raise the possibility that related topological features, including protected boundary modes, may be realizable in complex porous crystals.
@@ Appendix A Fourier Transformed Rigidity Matrix
$\mathbf{e}(\mathbf{k})=\mathbf{R}(\mathbf{k})\,\mathbf{u}(\mathbf{k})$ (13)
@@
$V=\tfrac{1}{2}\,\mathbf{u}^{\mathrm{T}}(R^{\mathrm{T}}KR)\mathbf{u},$ (14)
@@ Appendix B Code and Data Availability
All computational scripts and input files used in this work are available at:
Structures with CIF sizes above 35 kB were excluded to avoid memory bottlenecks. Neighbor environments were identified using CrystalNN[20] with a search cutoff of 25 Å. For the main dataset we used the default, chemically aligned settings with x_diff_weight = 1.5, which produce coordination numbers consistent with known metal and linker environments. For the UiO-66 tuning in Fig. 5, we instead reran CrystalNN with chemical heuristics effectively turned off (distance_cutoffs = None, weighted_cn = True, x_diff_weight = 0) and then retained only neighbors with $\mathrm{weight}\geq\tau$. In this case the network is controlled purely by geometry through the $\tau$ filter: bonds and angles present in both the default and $\tau$ filtered networks are treated as primary UFF4MOF constraints, while those that appear only in the $\tau$ filtered network are interpreted as auxiliary long range constraints and assigned uniform spring constants (see the repository for specific values). The Maxwell–Calladine index was consistently evaluated at the $\Gamma$-point, while phonon dispersions were sampled along Setyawan–Curtarolo [26] high-symmetry paths using 100 evenly spaced points per segment.
@@ Appendix C Silicon Benchmark
Since our present implementation builds the rigidity matrix from an explicit real-space reference cell containing a complete set of distinct bond and angle constraints, we compute the phonon dispersion of silicon using the standard eight-atom conventional cubic cell rather than the two-atom primitive cell. The comparison in Fig. 7 highlights several important features. The overall frequency range predicted by the rigidity model is higher than in the DFT reference calculation, reflecting the stiffer effective force constants in the harmonic spring network. Along the $\Gamma\rightarrow X$ path, however, the dispersions are in close agreement, with comparable band shapes and similar degeneracies at the high-symmetry points. Along the $X\rightarrow K$ path the two spectra begin to diverge, yet the qualitative structure is still preserved. The most noticeable differences appear along the $\Gamma\rightarrow L$ path, where the relative ordering and spacing of branches do not fully align with the DFT results. Despite these discrepancies, the simplified rigidity-based model reproduces the principal features of the silicon phonon dispersion with reasonable accuracy. This agreement supports the use of the approach for capturing the dominant degree of freedom structure in more complex materials.